A bank offers a savings account with the interest rate of 9% compounded daily. What is the effective interest rate? Use 360-day year.
Answer #1

Annual Interest Rate, APR = 9%
Number of compounding period per year, n = 360

Effective Annual Rate, EAR = (1 + APR/n)^n - 1
Effective Annual Rate, EAR = (1 + 0.09/360)^360 - 1
Effective Annual Rate, EAR = 1.00025^360 - 1
Effective Annual Rate, EAR = 1.0942 - 1
Effective Annual Rate, EAR = 0.0942 or 9.42%

So, effective interest rate is 9.42%